A wooden rectangular hollow box has outer edges of 6 inches and 5 inches, and height of 8 inches. The uniform thickness of the box is 1 inch.
a. find the amount of material used;
b. how many wooden blocks each cube with 1 inch edge will fit into the box?

Answer by ikleyn(52778)

.
Everything is clear to me in this post, except one thing: does the box have the upper lid or not ?

It seems that the problem must define it - but is does not.

From the other side, the problem does not mention about the top or bottom surfaces and says nothing about the base and so on . . .

Based on it, I will assume that the wooden material is at each of 6 faces of the box.

    Exterior dimensions of the block are 6, 5 and 8 inches.

    The volume of the box, calculated via exterior dimensions is  6*5*8 = 240 in^3.

    Interior dimensions are 6-2 = 4, 5-2 = 3 and 8-2 = 6 inches.

    The volume of the empty space inside the block, calculated via interior dimensions is  4*3*6 = 72 in^3.

    The volume of the material used = 240 - 72 = 168 in^3.

    How many wooden blocks . . . = 168.

I answered all questions.